# Optimal Fairness under Local Differential Privacy

### v1

@@ I Introduction
Machine learning algorithms are increasingly used in high-stakes decision-making tasks, making it essential to ensure fairness in their outcomes. Fairness in machine learning aims to prevent models from discriminating against demographic groups characterized by sensitive attributes such as gender or race. In practice, fairness is context-dependent, and no single definition applies universally. Metrics such as statistical parity [feldman2015certifying] and equalized opportunity [hardt2016equality] formalize different ways of assessing whether model decisions remain independent of sensitive attributes. Beyond fairness, privacy is another core component of trustworthy machine learning, as it protects individuals’ data in training datasets. Differential privacy (DP) [dwork2006calibrating, dwork2006differential] is the de facto standard for providing formal privacy guarantees in machine learning. DP is studied in both the central model, where a trusted curator is assumed, and the local DP (LDP) model, which assumes no trusted entity. Prior work shows that central DP can worsen fairness and may even be fundamentally incompatible with it [bagdasaryan2019differential, ganev2022robin, pujol2020fair, farrand2020neither, cummings2019compatibility]. In contrast, the interaction between fairness and LDP remains largely unexplored, and no analogous incompatibility result is known.
@@
Widely deployed systems using LDP algorithms have been implemented by companies such as Google, Apple, and Microsoft [erlingsson2014rappor, Apple_Privacy, ding2017collecting]. Although some work has examined the intersection of fairness and LDP, key challenges remain. For instance, [mozannar2020fair] proposes a learning scheme for training non-discriminatory classifiers with theoretical guarantees when only a privatized version of the sensitive attribute is available. Their approach employs an $\varepsilon$-LDP mechanism—generalized randomized response (GRR) [kairouz2014extremal]. Similarly, [chen2022fairness] considers a semi-private setting where a small subset of users reveal their sensitive attributes, while the rest apply an $\varepsilon$-LDP protocol. While these frameworks provide approaches for fair prediction with LDP-privatized data, they either focus exclusively on binary sensitive attributes or limit themselves to using GRR as the LDP mechanism, highlighting the need for further exploration.
@@
Recently, several works have begun examining the impact of LDP on fairness in classification [arcolezi2025group, makhlouf2024systematic, makhlouf2024impact]. For instance, [arcolezi2025group] empirically analyzed how different LDP mechanisms affect fairness and utility in binary classification, showing that privatizing sensitive attributes can improve fairness with minimal utility loss compared to training on non-private data. [makhlouf2024impact] showed that stronger privacy enhances fairness and that privatizing multiple sensitive attributes reduces unfairness more effectively than privatizing one. Complementing these empirical results, [makhlouf2024systematic] provided a theoretical analysis of how randomized response (RR) [warner1965randomized], the binary analogue of GRR, affects fairness under different privacy levels and data distributions. Their analysis identifies conditions under which privacy can either mitigate or exacerbate unfairness. Moreover, [xie2024privacy] showed that applying LDP during pre-processing limits sensitive information leakage and improves fairness.
@@
Fairness intervention methods are applied at different stages to mitigate bias: pre-processing before training [calders2009building, wang2019repairing, kamiran2012data, celis2020data, calmon2017optimized, hajian2012methodology, chakraborty2021bias, peng2022fairmask, gohar2023towards, madras2018learning, zemel2013learning, soen2023fair, yu2024fairbalance], in-processing during model training [lowy2021stochastic, cho2020fairdensity, cho2020fairmutual, jiang2020wasserstein, mary2019fairness, prost2019toward, zhang2018mitigating, agarwal2018reductions], and post-processing after generating predictions [wei2020optimized, wei2021optimized, chzhen2019leveraging, pleiss2017fairness, kim2020fact, jiang2020identifying, yang2020fairness, alghamdi2022beyond]. Among these, pre-processing methods offer the most flexibility within the data science pipeline, as they operate independently of the modeling algorithm [calmon2017optimized]. Our approach aligns closely with the pre-processing methods.
Building on recent evidence that LDP can promote fairness [arcolezi2025group, makhlouf2024systematic, makhlouf2024impact] and motivated by the flexibility of pre-processing methods, we study how to design the optimal LDP mechanism to reduce unfairness before model training. Specifically, we formulate the problem of identifying an LDP-based pre-processing mechanism that minimizes data unfairness. We further provide a theoretical analysis showing how reducing data unfairness translates to improved classification fairness. Finally, we validate the effectiveness of the optimal mechanisms through extensive experiments across diverse datasets and fairness metrics.
@@
  • We introduce the problem of designing LDP mechanisms that minimize data unfairness for both binary and non-binary sensitive attributes while preserving non-trivial utility. For the binary case, we derive a closed-form optimal solution, and for the non-binary case, we reformulate the problem as a min–max linear fractional program solvable via the branch-and-bound method [jiao2022solving].
@@ II Related Work
LDP and Fairness: Several studies have examined the impact of LDP on fairness in classification. [arcolezi2025group] empirically shows that applying LDP to sensitive attributes improves fairness with minimal utility loss compared to non-private data, highlighting that GRR and subset selection (SS) [wang2016mutual] offer the best accuracy-fairness-privacy trade-offs among state-of-the-art LDP mechanisms. Expanding on this, [makhlouf2024impact] experimentally demonstrates that stronger privacy (i.e., lower $\varepsilon$) further enhances fairness, with greater reductions in disparity when LDP is applied to multiple sensitive attributes. They subsequently delineate this observation by characterizing conditions under which RR implies lower unfairness for binary sensitive attributes [makhlouf2024systematic]. Other works focus on learning frameworks with privatized sensitive attributes. For instance, [mozannar2020fair] adapts non-discriminatory classifiers to work with privatized attributes using GRR, offering theoretical performance guarantees on utility and fairness. Similarly, [chen2022fairness] considers the scenario of semi-private sensitive attributes. Moreover, [xie2024privacy] shows that incorporating LDP randomizers during encoding can enhance fairness in representation learning while preserving utility. The work most closely related to ours is [makhlouf2024systematic], as they also theoretically analyze the impact of LDP on fairness. However, their study is limited to binary sensitive attributes and a specific LDP mechanism (RR) only.
@@
Central DP and Fairness: [bagdasaryan2019differential] empirically shows that central differential privacy, specifically differentially private stochastic gradient descent (DP-SGD) [abadi2016deep], has a disparate impact on accuracy of different subgroups. Similarly, [ganev2022robin] demonstrates that applying DP to synthetic data generation disproportionately affects minority sub-populations. [pujol2020fair] also finds that applying DP can exacerbate unfairness in decision-making tasks. [farrand2020neither] further shows that even with loose privacy guarantees and minimal data imbalance, DP-SGD can lead to disparate impact. [agarwal_2020] and [cummings2019compatibility] introduce an incompatibility result, highlighting that differential privacy and fairness are at odds when aiming for non-trivial accuracy in learning algorithms. Recently, [ko2024fairness] examined the impact of central DP on fairness beyond classification tasks. They demonstrated that using DP to set sampling rates for collecting socio-demographic data reduces unfairness across different population segments in resource allocation. Additionally, some works explore the intersection of privacy and fairness beyond the scope of central DP or group fairness. For instance, [chang2021privacy] examines the privacy risks of achieving group fairness, showing that fairness may compromise privacy, particularly through membership inference attacks, even when differential privacy is not explicitly applied. [dwork2012fairness] provides a theoretical link between individual fairness and DP, arguing that individual fairness can be seen as a generalization of DP and outlining conditions under which a DP mechanism ensures individual fairness.
@@
Pre-processing Techniques for Fairness: Pre-processing techniques for bias mitigation in fairness literature involve making changes to training data. Some methods modify values within the training data, such as altering ground truth labels (relabeling) [calders2009building, hajian2012methodology, kamiran2012data] or adjusting other features (perturbation) [wang2019repairing]. Methods such as [kamiran2012data, yu2024fairbalance] mitigate bias through reweighting of existing samples, while [wang2019repairing] perturbs the input distribution for disadvantaged groups to create a counterfactual distribution, particularly targeting binary sensitive attributes. Sampling methods adjust training data by changing sample distributions (e.g., adding or removing samples) or adapting their influence on training [chakraborty2021bias, celis2020data]. [celis2020data] proposes an optimization-based framework to learn distributions over the data domain that stay close to the empirical distribution. This technique is applicable to datasets with discrete or categorical attributes, both sensitive and non-sensitive. Other methods augment training data with additional, ideally unbiased features [madras2018learning]. Finally, some techniques learn transformations of the training data to reduce bias while retaining as much information as possible [zemel2013learning, calmon2017optimized]. [calmon2017optimized] introduces an optimization algorithm that modifies non-sensitive features and labels while keeping sensitive attributes unchanged, focusing on datasets with categorical or discrete attributes. Our method aligns with perturbation-based approaches, as we perturb the sensitive attribute using LDP.
@@ III-B Fairness Metrics
Definition 1 (Data unfairness $\Delta$[calmon2017optimized]).
@@
Definition 2 (Data unfairness $\Delta^{\prime}$ [kamiran2012data]).
@@
Definition 3 (Statistical parity [feldman2015certifying]).
@@
Definition 4 (Equalized opportunity [hardt2016equality]).
@@
Definition 5 ((Mean) equalized odds [hardt2016equality, alghamdi2022beyond]).
@@
It is worth noting that definitions of classification unfairness extend beyond the three mentioned here. Numerous definitions exist that may be useful to consider depending on the specific context of the problem [chouldechova2017fair, dwork2012fairness, berk2021fairness, corbett2017algorithmic, kilbertus2017avoiding, kleinberg2016inherent, kusner2017counterfactual, sabato2024fairnessunfairnessbinarymulticlass].
@@ III-C Discrimination-Accuracy Optimality
Definition 6 (Discrimination-accuracy optimal classifier [kamiran2012data]).
@@ III-D Local Differential Privacy
Definition 7 (Local differential privacy [warner1965randomized, evfimievski2003limiting]).
@@ III-E Common LDP Mechanisms
Generalized randomized response (GRR)[warner1965randomized, kairouz2014extremal]: Given a sensitive attribute $A$ taking values in $[k]$, the generalized randomized response (GRR) mechanism perturbs the true value $a\in[k]$ to preserve privacy. GRR outputs the true value $a$ with probability $\pi$, and any other possible value $a^{\prime}\in[k]\setminus\{a\}$ with a different probability $\bar{\pi}$. Specifically, for the perturbed output $Z$:
@@
Subset selection (SS) [wang2016mutual]: Given a sensitive attribute $A$ taking values in $[k]$, SS mechanism perturbs the true value $a\in[k]$ by reporting a subset of values $\Omega\subseteq[k]$. The mechanism aims to include the true value $a$ in the subset $\Omega$ with a higher probability than any other value in $k]\setminus\{a\}$. It is proved that the optimal subset size is $\omega=\left\lfloor\frac{k}{e^{\varepsilon}+1}\right\rceil$ [ye2018optimal]. The SS mechanism proceeds as follows:
@@ IV Main Results
As highlighted in Introduction, [arcolezi2023local] empirically demonstrates that applying different LDP mechanisms to the sensitive attributes of a dataset, followed by classifier training, can enhance classification fairness without significantly sacrificing model utility. This improvement is observed when compared to a non-private scenario, where the classifier is trained directly on the original non-private data in binary classification tasks. In our setting, sensitive attributes such as race and gender are privacy-sensitive. Therefore, we ensure privacy by applying LDP mechanisms to sensitive attributes, following the approach of [friedberg2023privacy], which guarantees LDP for each user’s group membership.
In this work, we aim to provide a theoretical foundation for this empirical observation. In Section IV-A, we begin by analyzing GRR as an LDP approach that offers a highly effective trade-off between fairness, utility, and privacy [arcolezi2023local]. Our goal is to understand how GRR impacts data fairness. In Section IV-B, we focus on the case of a binary sensitive attribute ($A\in\{0,1\}$) and identify the optimal LDP mechanism that can minimize unfairness in the dataset. Then, in Section IV-C, we extend our analysis to non-binary sensitive attributes ($A\in[k],k>2$). Here, we generalize the optimal LDP mechanism for fairness, which leads to solving a min-max linear fractional programming [jiao2022solving]. Finally, in Section IV-D, we provide theoretical conditions that establish the connection between data unfairness and classification unfairness.
@@ IV-C Optimal LDP Mechanism: Non-Binary Sensitive Attributes
where $q_{ij}$ represents the element in the $i$-th row and $j$-th column of the matrix $\mathbf{Q}$. Similar to Section IV-B, the perturbed dataset is $D_{M}=\{(x_{i},z_{i},y_{i})\}_{i=1}^{n}$. The idea of the perturbation matrix $\mathbf{Q}$ follows [celis2021fair], where this matrix is used to generate noisy sensitive attributes. However, in their setting, the perturbation mechanism is not an LDP mechanism. The matrix $\mathbf{Q}$ should satisfy the following constraints:
@@
Note that the parameters of this optimization problem are the non-diagonal entries of the matrix $\mathbf{Q}$, rather than all of its entries. The optimization problem (10) can be cast as a min–max linear fractional program [jiao2022solving], with an objective given by a ratio of two linear functions under linear inequality constraints. For fixed values of $\varepsilon$ and $\zeta$, this problem can be solved numerically using the branch-and-bound method, as outlined in [jiao2022solving]. Thus, for any given data distribution, we can numerically solve the problem to determine the optimal pre-processing mechanism with the desired $\varepsilon$ parameter for LDP and $\zeta$ error parameter. We choose Definition 1 over Definition 2 because the optimization problem (9) can be expressed as a min-max linear fractional program when using the data unfairness metric $\Delta$.

### v2

@@ I Introduction
Machine learning algorithms are increasingly used in high-stakes decision-making tasks, making it essential to ensure fairness in their outcomes. Fairness in machine learning aims to prevent models from discriminating against demographic groups characterized by sensitive attributes such as gender or race. In practice, fairness is context-dependent, and no single definition applies universally. Metrics such as statistical parity [32] and equalized opportunity [37] formalize different ways of assessing whether model decisions remain independent of sensitive attributes. Beyond fairness, privacy is another core component of trustworthy machine learning, as it protects individuals’ data in training datasets. Differential privacy (DP) [27, 28] is the de facto standard for providing formal privacy guarantees in machine learning. DP is studied in both the central model, where a trusted curator is assumed, and the local DP (LDP) model, which assumes no trusted entity. Prior work shows that central DP can worsen fairness and may even be fundamentally incompatible with it [8, 34, 58, 31, 22]. In contrast, the interaction between fairness and LDP remains largely unexplored, and no analogous incompatibility result is known.
@@
Widely deployed systems using LDP algorithms have been implemented by companies such as Google, Apple, and Microsoft [29, 23, 24]. Although some work has examined the intersection of fairness and LDP, key challenges remain. For instance, [54] proposes a learning scheme for training non-discriminatory classifiers with theoretical guarantees when only a privatized version of the sensitive attribute is available. Their approach employs an $\varepsilon$-LDP mechanism—generalized randomized response (GRR) [41]. Similarly, [16] considers a semi-private setting where a small subset of users reveal their sensitive attributes, while the rest apply an $\varepsilon$-LDP protocol. While these frameworks provide approaches for fair prediction with LDP-privatized data, they either focus exclusively on binary sensitive attributes or limit themselves to using GRR as the LDP mechanism, highlighting the need for further exploration.
@@
Recently, several works have begun examining the impact of LDP on fairness in classification [7, 52, 51]. For instance, [7] empirically analyzed how different LDP mechanisms affect fairness and utility in binary classification, showing that privatizing sensitive attributes can improve fairness with minimal utility loss compared to training on non-private data. [51] showed that stronger privacy enhances fairness and that privatizing multiple sensitive attributes reduces unfairness more effectively than privatizing one. Complementing these empirical results, [52] provided a theoretical analysis of how randomized response (RR) [65], the binary analogue of GRR, affects fairness under different privacy levels and data distributions. Their analysis identifies conditions under which privacy can either mitigate or exacerbate unfairness. Moreover, [69] showed that applying LDP during pre-processing limits sensitive information leakage and improves fairness.
@@
Fairness intervention methods are applied at different stages to mitigate bias: pre-processing before training [10, 63, 42, 13, 11, 36, 14, 55, 35, 50, 73, 60, 72], in-processing during model training [49, 17, 18, 39, 53, 57, 74, 2], and post-processing after generating predictions [66, 67, 20, 56, 45, 38, 70, 4]. Among these, pre-processing methods offer the most flexibility within the data science pipeline, as they operate independently of the modeling algorithm [11]. Our approach aligns closely with the pre-processing methods.
Building on recent evidence that LDP can promote fairness [7, 52, 51] and motivated by the flexibility of pre-processing methods, we study how to design the optimal LDP mechanism to reduce unfairness before model training. Specifically, we formulate the problem of identifying an LDP-based pre-processing mechanism that minimizes data unfairness. We further provide a theoretical analysis showing how reducing data unfairness translates to improved classification fairness. Finally, we validate the effectiveness of the optimal mechanisms through extensive experiments across diverse datasets and fairness metrics.
@@
  • We introduce the problem of designing LDP mechanisms that minimize data unfairness for both binary and non-binary sensitive attributes while preserving non-trivial utility. For the binary case, we derive a closed-form optimal solution, and for the non-binary case, we reformulate the problem as a min–max linear fractional program solvable via the branch-and-bound method [40].
@@ II Related Work
LDP and Fairness: Several studies have examined the impact of LDP on fairness in classification. [7] empirically shows that applying LDP to sensitive attributes improves fairness with minimal utility loss compared to non-private data, highlighting that GRR and subset selection (SS) [64] offer the best accuracy-fairness-privacy trade-offs among state-of-the-art LDP mechanisms. Expanding on this, [51] experimentally demonstrates that stronger privacy (i.e., lower $\varepsilon$) further enhances fairness, with greater reductions in disparity when LDP is applied to multiple sensitive attributes. They subsequently delineate this observation by characterizing conditions under which RR implies lower unfairness for binary sensitive attributes [52]. Other works focus on learning frameworks with privatized sensitive attributes. For instance, [54] adapts non-discriminatory classifiers to work with privatized attributes using GRR, offering theoretical performance guarantees on utility and fairness. Similarly, [16] considers the scenario of semi-private sensitive attributes. Moreover, [69] shows that incorporating LDP randomizers during encoding can enhance fairness in representation learning while preserving utility. The work most closely related to ours is [52], as they also theoretically analyze the impact of LDP on fairness. However, their study is limited to binary sensitive attributes and a specific LDP mechanism (RR) only.
@@
Central DP and Fairness: [8] empirically shows that central differential privacy, specifically differentially private stochastic gradient descent (DP-SGD) [1], has a disparate impact on accuracy of different subgroups. Similarly, [34] demonstrates that applying DP to synthetic data generation disproportionately affects minority sub-populations. [58] also finds that applying DP can exacerbate unfairness in decision-making tasks. [31] further shows that even with loose privacy guarantees and minimal data imbalance, DP-SGD can lead to disparate impact. [3] and [22] introduce an incompatibility result, highlighting that differential privacy and fairness are at odds when aiming for non-trivial accuracy in learning algorithms. Recently, [47] examined the impact of central DP on fairness beyond classification tasks. They demonstrated that using DP to set sampling rates for collecting socio-demographic data reduces unfairness across different population segments in resource allocation. Additionally, some works explore the intersection of privacy and fairness beyond the scope of central DP or group fairness. For instance, [15] examines the privacy risks of achieving group fairness, showing that fairness may compromise privacy, particularly through membership inference attacks, even when differential privacy is not explicitly applied. [26] provides a theoretical link between individual fairness and DP, arguing that individual fairness can be seen as a generalization of DP and outlining conditions under which a DP mechanism ensures individual fairness.
@@
Pre-processing Techniques for Fairness: Pre-processing techniques for bias mitigation in fairness literature involve making changes to training data. Some methods modify values within the training data, such as altering ground truth labels (relabeling) [10, 36, 42] or adjusting other features (perturbation) [63]. Methods such as [42, 72] mitigate bias through reweighting of existing samples, while [63] perturbs the input distribution for disadvantaged groups to create a counterfactual distribution, particularly targeting binary sensitive attributes. Sampling methods adjust training data by changing sample distributions (e.g., adding or removing samples) or adapting their influence on training [14, 13]. [13] proposes an optimization-based framework to learn distributions over the data domain that stay close to the empirical distribution. This technique is applicable to datasets with discrete or categorical attributes, both sensitive and non-sensitive. Other methods augment training data with additional, ideally unbiased features [50]. Finally, some techniques learn transformations of the training data to reduce bias while retaining as much information as possible [73, 11]. [11] introduces an optimization algorithm that modifies non-sensitive features and labels while keeping sensitive attributes unchanged, focusing on datasets with categorical or discrete attributes. Our method aligns with perturbation-based approaches, as we perturb the sensitive attribute using LDP.
@@ III-B Fairness Metrics
Definition 1 (Data unfairness $\Delta$[11]).
@@
Definition 2 (Data unfairness $\Delta^{\prime}$ [42]).
@@
Definition 3 (Statistical parity [32]).
@@
Definition 4 (Equalized opportunity [37]).
@@
Definition 5 ((Mean) equalized odds [37, 4]).
@@
It is worth noting that definitions of classification unfairness extend beyond the three mentioned here. Numerous definitions exist that may be useful to consider depending on the specific context of the problem [19, 26, 9, 21, 44, 46, 48, 59].
@@ III-C Discrimination-Accuracy Optimality
Definition 6 (Discrimination-accuracy optimal classifier [42]).
@@ III-D Local Differential Privacy
Definition 7 (Local differential privacy [65, 30]).
@@ III-E Common LDP Mechanisms
Generalized randomized response (GRR)[65, 41]: Given a sensitive attribute $A$ taking values in $[k]$, the generalized randomized response (GRR) mechanism perturbs the true value $a\in[k]$ to preserve privacy. GRR outputs the true value $a$ with probability $\pi$, and any other possible value $a^{\prime}\in[k]\setminus\{a\}$ with a different probability $\bar{\pi}$. Specifically, for the perturbed output $Z$:
@@
Subset selection (SS) [64]: Given a sensitive attribute $A$ taking values in $[k]$, SS mechanism perturbs the true value $a\in[k]$ by reporting a subset of values $\Omega\subseteq[k]$. The mechanism aims to include the true value $a$ in the subset $\Omega$ with a higher probability than any other value in $k]\setminus\{a\}$. It is proved that the optimal subset size is $\omega=\max\Big(1,\bigl\lfloor\tfrac{k}{e^{\varepsilon}+1}\bigr\rceil\Big)$ [71]. The SS mechanism proceeds as follows:
@@ IV Main Results
As highlighted in Introduction, [6] empirically demonstrates that applying different LDP mechanisms to the sensitive attributes of a dataset, followed by classifier training, can enhance classification fairness without significantly sacrificing model utility. This improvement is observed when compared to a non-private scenario, where the classifier is trained directly on the original non-private data in binary classification tasks. In our setting, sensitive attributes such as race and gender are privacy-sensitive. Therefore, we ensure privacy by applying LDP mechanisms to sensitive attributes, following the approach of [33], which guarantees LDP for each user’s group membership.
In this work, we aim to provide a theoretical foundation for this empirical observation. In Section IV-A, we begin by analyzing GRR as an LDP approach that offers a highly effective trade-off between fairness, utility, and privacy [6]. Our goal is to understand how GRR impacts data fairness. In Section IV-B, we focus on the case of a binary sensitive attribute ($A\in\{0,1\}$) and identify the optimal LDP mechanism that can minimize unfairness in the dataset. Then, in Section IV-C, we extend our analysis to non-binary sensitive attributes ($A\in[k],k>2$). Here, we generalize the optimal LDP mechanism for fairness, which leads to solving a min-max linear fractional programming [40]. Finally, in Section IV-D, we provide theoretical conditions that establish the connection between data unfairness and classification unfairness.
@@ IV-C Optimal LDP Mechanism: Non-Binary Sensitive Attributes
where $q_{ij}$ represents the element in the $i$-th row and $j$-th column of the matrix $\mathbf{Q}$. Similar to Section IV-B, the perturbed dataset is $D_{M}=\{(x_{i},z_{i},y_{i})\}_{i=1}^{n}$. The idea of the perturbation matrix $\mathbf{Q}$ follows [12], where this matrix is used to generate noisy sensitive attributes. However, in their setting, the perturbation mechanism is not an LDP mechanism. The matrix $\mathbf{Q}$ should satisfy the following constraints:
@@
Note that the parameters of this optimization problem are the non-diagonal entries of the matrix $\mathbf{Q}$, rather than all of its entries. The optimization problem (10) can be cast as a min–max linear fractional program [40], with an objective given by a ratio of two linear functions under linear inequality constraints. For fixed values of $\varepsilon$ and $\zeta$, this problem can be solved numerically using the branch-and-bound method, as outlined in [40]. Thus, for any given data distribution, we can numerically solve the problem to determine the optimal pre-processing mechanism with the desired $\varepsilon$ parameter for LDP and $\zeta$ error parameter. We choose Definition 1 over Definition 2 because the optimization problem (9) can be expressed as a min-max linear fractional program when using the data unfairness metric $\Delta$.